# Model Supply Chain Poisoning: Backdooring Pre-trained Models via Embedding Indistinguishability

### v2

@@ Abstract.
Backdoor attack, Pre-trained model, Model supply chain
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the ACM Web Conference 2025; April 28 – May 02, 2025; Sydney, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 02, 2025, Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06submissionid: 1525ccs: Computing methodologies Computer vision
@@ 1. Introduction
Task-agnostic backdoor attacks have emerged to address these constraints (Jia et al., 2022; Zhang et al., 2023; Tao et al., 2024; Shen et al., 2021). Due to the absence of downstream labels, the key to achieve task-agnostic backdoor attacks lies in aligning a backdoor trigger with the pre-defined reference embedding. Such embedding can hit a target label of various downstream tasks after fine-tuning. The reference embedding is usually predicted from the target class images (shadow images) (Jia et al., 2022; Tao et al., 2024) or created manually using empirical methods (Zhang et al., 2023; Shen et al., 2021). The manually created reference embeddings are also known as artificially pre-defined output representations (PORs).
We conduct extensive evaluations of TransTroj on four widely used PTMs – ResNet, VGG, ViT, and CLIP – and six downstream tasks, including CIFAR-10, CIFAR-100, GTSRB, Caltech 101, Caltech 256, and Oxford-IIIT Pet. Our experimental results demonstrate that TransTroj significantly outperforms SOTA task-agnostic backdoor attacks, achieving nearly 100% attack success rates on most downstream tasks with minimal impact on model accuracy. Moreover, TransTroj exhibits robustness under various system settings and remains effective even when subjected to model reconstruction-based defenses.
Our key contributions are summarized as follows:
  • We propose TransTroj, a novel backdoor attack that is functionality-preserving, durable, and truly task-agnostic, effectively transferring through the model supply chain.
  • We design a two-stage optimization framework that separately optimizes the trigger and the victim PTM to achieve embedding indistinguishability without sacrificing model performance on clean data.
  • We provide comprehensive experimental evidence demonstrating the effectiveness and robustness of TransTroj across multiple PTMs and downstream tasks, achieving nearly 100% attack success rates on most downstream tasks.
Figure 1. Illustration of transferable backdoor attacks. The adversary injects backdoor into a clean PTM and launches attack when the backdoored PTM is leveraged to fine-tune downstream tasks. Note that the task-agnostic backdoor can be activated in various downstream tasks.
2. Related Work and Comparisons
Table 1. Comparison of backdoor attack methods on pre-trained models. Prior knowledge: the attack requires specific downstream task knowledge; Task-agnostic: backdoors can be activated across various downstream tasks; Durable: the attack maintains a high attack success rate even after fine-tuning.
@@ 2. Related Work and Comparisons
Task-specific backdoor attacks. Existing task-specific backdoor attacks (Yao et al., 2019; Kurita et al., 2020; Li et al., 2021; Zhang et al., 2021) against PTMs usually require prior knowledge of the downstream tasks. For example, Kurita et al. (Kurita et al., 2020) proposed RIPPLe, which directly poisons the fine-tuned downstream model and then acquires the PTM part as a task-specific backdoored PTM. Yao et al. (Yao et al., 2019) proposed a latent backdoor attack (LBA) to inject backdoors into a teacher classifier built on PTMs. To successfully attack the specific downstream task, LBA also needs a labeled dataset that is similar to the target downstream dataset. When the backdoored teacher classifier is used to fine-tune a student classifier for the specific target downstream task, the student classifier inherits the backdoor behavior. Other backdoor attacks are designed for specific pre-training paradigms like contrastive learning (Carlini and Terzis, 2022; Saha et al., 2020; Zhang et al., 2024), masked image modeling (Shen et al., 2022), and masked language modeling (Chen et al., 2021; Du et al., 2023; Shen et al., 2022).
Task-agnostic backdoor attacks. Compared to task-specific attacks, task-agnostic attacks (Zhang et al., 2023; Jia et al., 2022; Tao et al., 2024) are more severe threats as they can compromise various downstream tasks built on the PTMs. BadEncoder(Jia et al., 2022) exemplified a backdoor attack on self-supervised learning, which compromises the downstream classifier by inserting backdoors into an image encoder. It achieves a high attack success rate in transfer learning scenarios without fine-tuning PTMs, such as linear probing and zero-shot classification. NeuBA (Zhang et al., 2023) trained PTMs to build strong links between triggers and manually pre-defined output representations (PORs). After fine-tuning, the POR can hit a certain label of the downstream task. In binary classification tasks, the PORs of NueBA can cover positive and negative classes and accomplish targeted attacks. Unfortunately, existing attacks are unable to maintain a high attack success rate after fine-tuning or be fully task-agnostic. To the best of our knowledge, TransTroj is the first method that simultaneously satisfies functionality-preserving, durable, and task-agnostic.
@@ 3.2. Threat Model
Adversary’s goal. In our threat model, the model publisher $\mathbf{MP}$ is malicious, where he/her trains and releases an evil PTM, denoted as $\tilde{F}$, with a “transferable” backdoor. The goal is that any downstream model $\tilde{f}$ developed based on $\tilde{F}$ will inherit the backdoor. As shown in Fig. 1, the adversary initially prepares a clean PTM $F$. He/she selects a target classes $y_{t}$ and optimizes the corresponding trigger $\boldsymbol{t}$. Then, he/she crafts a backdoored PTM $\tilde{F}$ to bind the trigger $\boldsymbol{t}$ with the target class $y_{t}$. When a victim $\mathbf{DU}$ downloads the backdoored PTM $\tilde{F}$ and adapts it to his/her downstream tasks through fine-tuning, the adversary can activate the backdoor in the downstream model by querying with samples that contains the trigger $\boldsymbol{t}$. A transferable backdoor should meet the following goals:
@@
    Durable. Fine-tuning the PTM weights can lead the model to forget some previously learned knowledge, a phenomenon known as catastrophic forgetting (McCloskey and Cohen, 1989). A transferable backdoor must resist being forgotten during the fine-tuning process.
@@ 4.3. Pre- and Post-indistinguishability
On other hand, post-indistinguishability pertains to the indistinguishability of embeddings from a backdoored PTM, whcih is formally defined as follows.
@@ 4.4. Two-Stage Optimization
Thus, the backdoored PTM can be derived from this optimization problem:
@@ 5.1. Experimental Setup
To extensively evaluate our TransTroj, we conduct an series of experiments across a diverse range of pre-trained models and downstream tasks.
@@ 5.5. Sensitivity Analysis
We investigate various factors, such as trigger infinity norm and shadow dataset size, that may impact the performance of TransTroj. For these studies, ResNet-18 is chosen as the PTM, Caltech 256 serves as the downstream task, and “Sunflower” is selected as the target class.
Trigger infinity norm. Gradually reduce the infinity norm $\xi$ of triggers from 10 to 3, we evaluate the attack performance on Caltech 256, as shown in Fig. 8 (left). It’s obvious that the attack success rate significantly decreases with the reduction of the infinity norm, due to the failure of trigger optimization to make the poisoned embeddings resemble the clean embeddings when the infinity norm is excessively small. In other words, pre-indistinguishability cannot be ensured.
@@
Ablation study. We conduct a comprehensive ablation study to evaluate the effectiveness of our method, including experiments on the impact of different loss terms, trigger patterns, and the use of reference images. For more detailed analysis and results, please refer to Appendix C.
@@ Appendix C Ablation Study
Loss terms. Our TransTroj incorporates three loss terms, i.e.$\mathcal{L}_{\mathrm{pre}}$, $\mathcal{L}_{\mathrm{post}}$ and $\mathcal{L}_{\mathrm{func}}$. We also include a hyperparameter $\lambda$ to balance $\mathcal{L}_{\mathrm{post}}$ and $\mathcal{L}_{\mathrm{func}}$. Therefore, it is important to examine the impact of $\lambda$ on our method. The results of this examination are displayed in Fig. 13. We observe that both an exceedingly large or samll $\lambda$ prevents TransTroj from achieving a high attack success rate while preserving accuracy. In particular, the model accuracy starts to decrease when $\lambda$ exceeds approximately $10$. Therefore, both $\mathcal{L}_{\mathrm{post}}$ and $\mathcal{L}_{\mathrm{func}}$ are important for effective backdoor training.

### v3

@@ Abstract.
Backdoor attack; Pre-trained model; Model supply chain
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714624isbn: 979-8-4007-1274-6/25/04ccs: Security and privacy
@@ 1. Introduction
Table 1. Comparison of backdoor attacks on PTMs. Prior knowledge: the attack requires specific task knowledge; Task-agnostic: backdoors can be activated across various tasks; Durable: the attack maintains a high ASR even after fine-tuning.
@@
Task-agnostic backdoor attacks have emerged to address these constraints (Jia et al., 2022; Zhang et al., 2023; Tao et al., 2024; Shen et al., 2021). Due to the absence of downstream labels, the key to achieve task-agnostic backdoor attacks lies in aligning a backdoor trigger with the pre-defined reference embedding. Such embedding can hit a target label of various downstream tasks after fine-tuning. The reference embedding is usually predicted from the target class images (shadow images) (Jia et al., 2022; Tao et al., 2024) or created manually (Zhang et al., 2023; Shen et al., 2021). The manually created reference embeddings are also known as artificially pre-defined output representations (PORs).
We conduct extensive evaluations of TransTroj on four PTMs and six downstream tasks. Our experimental results demonstrate that TransTroj significantly outperforms SOTA task-agnostic backdoor attacks, achieving nearly 100% attack success rates on most tasks with minimal impact on model accuracy. Moreover, TransTroj exhibits robustness under various system settings and remains effective even when subjected to model reconstruction-based defenses.
Our key contributions are summarized as follows:
  • We propose TransTroj, a novel backdoor attack which is functionality-preserving, durable, and truly task-agnostic, effectively transferring through the model supply chain.
  • We design a two-stage optimization framework that separately optimizes the trigger and the victim PTM to achieve embedding indistinguishability.
  • We provide comprehensive experimental evidence demonstrating the effectiveness and robustness of TransTroj across multiple PTMs and downstream tasks.
2. Related Work and Comparisons
@@
Task-specific backdoor attacks. Existing task-specific backdoor attacks (Yao et al., 2019; Kurita et al., 2020; Li et al., 2021; Zhang et al., 2021) against PTMs usually require prior knowledge of the downstream tasks. For example, Kurita et al. (Kurita et al., 2020) proposed RIPPLe, which directly poisons the fine-tuned downstream model and then acquires the PTM part as a task-specific backdoored PTM. Yao et al. (Yao et al., 2019) proposed a latent backdoor attack (LBA) to inject backdoors into a teacher classifier built on PTMs. To successfully attack the specific downstream task, LBA also needs a labeled dataset that is similar to the target downstream dataset. When the backdoored teacher classifier is used to fine-tune a student classifier for the specific target downstream task, the student classifier inherits the backdoor behavior. Other backdoor attacks are designed for specific pre-training paradigms like contrastive learning (Carlini and Terzis, 2022; Saha et al., 2020; Zhang et al., 2024) and masked language modeling (Chen et al., 2021; Du et al., 2023; Shen et al., 2022).
Task-agnostic backdoor attacks. Compared to task-specific attacks, task-agnostic attacks (Zhang et al., 2023; Jia et al., 2022; Tao et al., 2024) are more severe threats as they can compromise various downstream tasks built on the PTMs. BadEncoder(Jia et al., 2022) exemplified a backdoor attack on self-supervised learning, which compromises the downstream classifier by inserting backdoors into an image encoder. It achieves a high attack success rate in transfer learning scenarios without fine-tuning PTMs.NeuBA (Zhang et al., 2023) trained PTMs to build strong links between triggers and manually pre-defined output representations (PORs). After fine-tuning, the POR can hit a certain label of the downstream task. In binary classification tasks, the PORs of NueBA can cover positive and negative classes and accomplish targeted attacks. Unfortunately, existing attacks are unable to maintain a high attack success rate after fine-tuning or be fully task-agnostic. To the best of our knowledge, TransTroj is the first method that simultaneously satisfies functionality-preserving, durable, and task-agnostic.
@@ 3.1. System Model
Figure 1. Illustration of transferable backdoor attacks. The adversary injects backdoor into a clean PTM and launches attack when the backdoored PTM is leveraged to fine-tune downstream tasks.
@@ 3.2. Threat Model
Adversary’s goal. In our threat model, the model publisher $\mathbf{MP}$ is malicious, where he/she trains and releases an evil PTM, denoted as $\tilde{F}$, with a “transferable” backdoor. The goal is that any downstream model $\tilde{f}$ developed based on $\tilde{F}$ will inherit the backdoor. As shown in Fig. 1, the adversary initially prepares a clean PTM $F$. He/she selects a target classes $y_{t}$ and optimizes the corresponding trigger $\boldsymbol{t}$. Then, he/she crafts a backdoored PTM $\tilde{F}$ to bind the trigger $\boldsymbol{t}$ with the target class $y_{t}$. When a victim $\mathbf{DU}$ downloads the backdoored PTM $\tilde{F}$ and adapts it to his/her downstream tasks through fine-tuning, the adversary can activate the backdoor in the downstream model by querying with samples that contains the trigger $\boldsymbol{t}$. A transferable backdoor should meet the following goals:
@@
    Durable. Fine-tuning the PTMs can lead the model to forget some previously learned knowledge, a phenomenon known as catastrophic forgetting (McCloskey and Cohen, 1989). A transferable backdoor must resist being forgotten during the fine-tuning process.
@@ 4.3. Pre- and Post-indistinguishability
On other hand, post-indistinguishability pertains to the indistinguishability of embeddings from a backdoored PTM, which is formally defined as follows.
@@ 4.4. Two-Stage Optimization
Thus, the backdoored PTM can be derived from the following optimization problem:
@@ 5.1. Experimental Setup
To extensively evaluate our TransTroj, we conduct a series of experiments across a diverse range of PTMs and downstream tasks.
@@ 5.5. Sensitivity Analysis
We analyze factors like trigger infinity norm and shadow dataset size affecting TransTroj’s performance, using ResNet-18 as the PTM, Caltech 256 as the task, and ”Sunflower” as the target class.
Trigger infinity norm. Gradually reduce the infinity norm $\xi$ of triggers from 10 to 3, we evaluate the attack performance on Caltech 256, as shown in Fig. 8 (left). It’s obvious that the ASR significantly decreases with the reduction of the infinity norm, due to the failure of trigger optimization to make the poisoned embeddings resemble the clean embeddings when the infinity norm is excessively small. In other words, pre-indistinguishability cannot be ensured.
@@
Ablation study. We carry out an extensive ablation study to assess the efficacy of our method. For more detailed analysis and results, please refer to Appendix C.
@@ 6. Conclusion
Acknowledgements.
This work was supported in part by the National Key R&D Program of China under Grant No. 2022YFB3103500; the National Natural Science Foundation of China under Grant No. 62472057; the Open Subject Fund of Key Laboratory of Cyberspace Security Defense under Grant No. 2024-MS-10; the National Research Foundation, Singapore and DSO National Laboratories under its AI Singapore Programme (AISG Award No. AISG2- GC-2023-008).
@@ Appendix C Ablation Study
Loss terms. Our TransTroj incorporates three loss terms, i.e.$\mathcal{L}_{\mathrm{pre}}$, $\mathcal{L}_{\mathrm{post}}$ and $\mathcal{L}_{\mathrm{func}}$. We also include a hyperparameter $\lambda$ to balance $\mathcal{L}_{\mathrm{post}}$ and $\mathcal{L}_{\mathrm{func}}$. Therefore, it is important to examine the impact of $\lambda$ on our method. The results of this examination are displayed in Fig. 13. We observe that both an exceedingly large or small $\lambda$ prevents TransTroj from achieving a high attack success rate while preserving accuracy. In particular, the model accuracy starts to decrease when $\lambda$ exceeds approximately $10$. Therefore, both $\mathcal{L}_{\mathrm{post}}$ and $\mathcal{L}_{\mathrm{func}}$ are important for effective backdoor training.